OpenStudy (sparklestaraa):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences

OpenStudy (dtan5457):

this is the same thing as the slope intercept form we did last time except your gonna replace the y with

OpenStudy (dtan5457):

f(x)

OpenStudy (sparklestaraa):

so how do i go about writing this? @dtan5457

OpenStudy (dtan5457):

your old equation after diving everything by 3 etc was y=-2/3x+400

OpenStudy (dtan5457):

now it's just f(x)=-2/3x+400

OpenStudy (sparklestaraa):

so how do i explain what the graph of the function represents

OpenStudy (dtan5457):

what do you think it represents?

OpenStudy (sparklestaraa):

the graph maybe represents the input of the graph

OpenStudy (dtan5457):

it represents everything "The profit on each smart phone case is $2, and the profit on each tablet case is $3."

OpenStudy (sparklestaraa):

ohhh ok :D that makes sense then

OpenStudy (sparklestaraa):

how do i Graph the function and on the graph, I have to make sure to label the intercepts...

OpenStudy (dtan5457):

is this on paper again

OpenStudy (sparklestaraa):

either paper or on a graph program online

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (dtan5457):

https://www.desmos.com/calculator

OpenStudy (dtan5457):

type it in and you get a nice picture

OpenStudy (sparklestaraa):

:D thanks soooo much

OpenStudy (sparklestaraa):

i did that it shows nothing @dtan5457

OpenStudy (dtan5457):

what are you suppose to type in again?

OpenStudy (sparklestaraa):

Graph the function and on the graph, I have to make sure to label the intercepts

OpenStudy (dtan5457):

so if it's online graphing, your allowed to take a screen shot?

OpenStudy (sparklestaraa):

yes :P

OpenStudy (dtan5457):

the function was y=-2/3x+400 right?

OpenStudy (sparklestaraa):

yes :D

OpenStudy (dtan5457):

send your teacher this

OpenStudy (sparklestaraa):

omg thanks soooo much :D

OpenStudy (sparklestaraa):

I still have to expand on this topic with other questions -,-

OpenStudy (dtan5457):

make a new question so others can help you along with me.

OpenStudy (sparklestaraa):

:D ok i will prob make it in the morning as it is late here